# Edge Universality for Inhomogeneous Random Matrices

### v1

@@ 2.1 Ribbon graph expansion
The correspondence between ribbon graphs on orientable surfaces and Hermitian random matrix ensembles (e.g., GUE) is classical and well-studied; see, for example, [HZ86, Kon92, Oko00, LZ04]. For real ensembles (e.g., GOE), similar but subtler relations exist, involving ribbon graphs on possibly non-orientable surfaces, as formalized via the notion of Möbius graphs in [MW03].
Informally, a ribbon graph is a graph embedded into a compact surface (possibly with boundary), such that the complement of the embedding is a disjoint union of open disks, called faces. It therefore defines a cell decomposition of the surface and uniquely determines its topological type (orientable or not). Recent developments include the study of ribbon graphs with boundary, which arise naturally in open intersection theory; see, e.g., [BT17, Tes23].
@@
Label the vertices and edges in cyclic order as $v_{1},\dots,v_{k}$, and define directed edges as $\vec{e}_{j}=\overrightarrow{v_{j}v_{\gamma(j)}}$.
@@
  • In the real case ($\beta=1$), glue either in the same or opposite direction (possibly yielding a non-orientable surface).
After gluing, we identify $v_{s}=v_{\gamma(t)}$ and $v_{t}=v_{\gamma(s)}$. Edges in $J^{c}$ are left unglued and correspond to the external deformation $A_{N}$, which are called open edges. The result is an $s$-cell ribbon graph with perimeter $(k_{1},\dots,k_{s})$ on a compact surface. For an illustration, see Figures 2 and 2.
@@
A diagram $\Gamma$ is a ribbon graph satisfying the following conditions:
@@
We introduce the following notation:
  • Let $E(\Gamma)$ denote the set of all edges of $\Gamma$, with $E_{\mathrm{int}}(\Gamma)$ and $E_{\mathrm{b}}(\Gamma)$ denoting the sets of interior and open (boundary) edges, respectively;
  • Let $V(\Gamma)$, $V_{\mathrm{int}}(\Gamma)$, and $V_{\mathrm{b}}(\Gamma)$ denote the sets of all, interior, and boundary vertices, respectively;
  • For each $j\in[s]$, let $\partial D_{j}\subset E(\Gamma)$ denote the set of edges forming the boundary of face $j$.
Let $\mathscr{D}_{s;\beta}$ be the set of all $s$-cell diagrams (on orientable surfaces when $\beta=2$). We write $\mathscr{D}_{s;\beta}^{*}\subset\mathscr{D}_{s;\beta}$ for the subset of connected diagrams.
@@
    Record a weight $w_{e}$ on each edge $e$ equal to one plus the number of divalent vertices removed along it.
@@
We now state the ribbon graph expansion for the deformed Gaussian IRM ensemble as established in [GLZ24].
Proposition 2.4 ([GLZ24]).
@@
with the convention that for each trivial diagram (i.e., a single isolated vertex), the contribution is $N\delta_{l_{i},0}$. Here, the outer sum is taken over all vertex labelings $\eta:V(\Gamma)\to[N]$.
Then for any integers $m_{1},\dots,m_{s}\geq 0$,
@@
This follows directly from Propositions 2.5 and 2.6 of [GLZ24]. ∎
@@ 2.2 Graphical expansion for Chebyshev moments
The Chebyshev polynomials of first and second kind $T_{n}(x)$ and $U_{n}(x)$ with $n\geq 0$ are defined as ,
@@
where the prime notation $\sum^{\prime}$ is over all $j\equiv m\pmod{2}$ and the term of $j=0$ is to be halved if there is one; see [MH03, eq (2.14)]. In the reverse direction the exact formula for Chebyshev polynomials in terms of powers of $x$ is
@@
Let $X$ be the deformed Gaussian matrix in Definition 1.1, then for any non-negative integers $n_{j}\geq 0$, there exists a diagrammatic expansion of the form
@@
[…]
@@ 3 Analysis of diagram functions
In this section, we focus on analyzing key analytic properties of diagram functions $F_{\Gamma}(\{n_{j}\})$ for any connected diagram $\Gamma\in\mathscr{D}_{s;\beta}^{*}$, including upper-bound estimates and precise asymptotic behavior.
@@ 3.1 Upper bound estimates
$\sum_{w_{e}\leq t_{N},e\in S\cap\mathcal{F}^{c}}\prod_{e\in S\cap\mathcal{F}^{c}}p_{w_{e}}(\eta(x),\eta(y))\leq\big{(}\frac{\gamma t_{N}}{N}\big{)}^{|S\cap\mathcal{F}^{c}|}.$ (3.7)
While for every edge from $S^{c}\cap\mathcal{F}^{c}$ we see from the assumption (B2) that $p_{w_{e}}(\eta(x),\eta(y))\leq(1+\delta)N^{-1}$, which implies
@@
This upper bound is irrelevant to the labeling on open vertices and weight on open edges. Since the open edges are also connected end to end into the union of several circles (as the boundary of Riemann surface are union of circles) for any given circle with $k$ open vertices, by the spectral decomposition $A_{N}=\sum_{t=1}^{r}a_{t}\mathbf{q}_{t}\mathbf{q}_{t}^{*}$ with $Q=(\mathbf{q}_{1},\ldots,\mathbf{q}_{N})$ defined in (1.5) we have
@@
If $|E|\geq\frac{n}{t_{N}}$, we also fix the labeling of one vertex and designate the remaining vertices as $V_{\mathrm{int}}$ with corresponding edges $E_{\mathrm{int}}$. Noting that $|V_{\mathrm{b}}|=|E_{\mathrm{b}}|=0$ and setting $a=0$, we follow the procedure in (3.18) to obtain
@@ 3.2 Asymptotics for diagram functions
$\bigg{|}\prod_{(x,y)\in E_{\mathrm{int}}}p_{w_{e}}(\eta(x),\eta(y))-\prod_{(x,y)\in E_{\mathrm{int}}}\frac{1}{N}\bigg{|}\mathbf{1}\left(\{w_{e}\geq t_{N}\}_{e\in S}\right)\cdot\mathbf{1}\left(\{w_{e}\geq t_{N}\}_{e\in S^{c}}\right).$ (3.24)
@@
Since $r,|E_{\mathrm{int}}|$ are fixed, the desired result
@@ 4.1 Mixed moments in Gaussian case
The construction of $\phi$ follows the vertex-splitting procedure introduced in [FS10] (see e.g. Step (iii) in Definition II.1.3, Figure 5 in [GLZ24] or Lemma 2.22 in [LZ24b]). The key idea is to iteratively replace each unmarked vertex of degree $d>3$ with a tree of trivalent vertices connected via auxiliary edges, such that the total number of edges increases by d3d-3 and the total number of vertices increases by d3d-3. Crucially, such a replacement preserves the difference |E||V||E|-|V|.
@@
Lemma 4.3 ([FS10, Proposition II.2.3]; [GLZ24, Proposition 3.7]).
@@
For the deformed Gaussian matrix $X$ in Definition 1.1 and under the assumptions (B1)-(B2) of Definition 1.2, with $n=\sum_{j=1}^{s}n_{j}$, $a=\|A_{N}\|_{\mathrm{op}}$, the following holds.
@@
    $\mathbb{E}\left[\prod_{j=1}^{s}\operatorname{Tr}\left(U_{n_{j}}\left(\frac{X}{2}\right)\right)\right]\leq(1+a^{n})(Cn)^{s}\left\{e^{C\frac{n^{\frac{3}{2}}}{\sqrt{N}}}+(1\lor\frac{t_{N}}{n})^{s}e^{C\frac{n^{2}t_{N}}{N}}+\Big{(}(\frac{Cn^{2}t_{N}}{N})^{[\frac{n}{t_{N}}]}\land 1\Big{)}e^{C\frac{n^{3}}{N}}\right\},$ (4.9)
@@
$|E|=3\ell+s,\quad|V|=2\ell+s,\quad|E_{\mathrm{int}}|-|V_{\mathrm{int}}|=|E|-|V|=\ell.$ (4.15)
Here we use the fact $|E_{\mathrm{b}}|=|V_{\mathrm{b}}|$, see [GLZ24, Lemma 2.9]. Rewriting $G_{\Gamma}(n):=G(n,\ell)$ that depends only on $\ell$ and $s$, we use Lemma 4.3 to further arrive at
@@
    It also worthy to note
@@
Note that $s$ is a fixed positive integer, put (4.18)-(4.22) together and we obtain
@@
is absolutely summable. Therefore, for any sufficiently large constant $C>0$, the contribution from those diagram functions $F_{\Gamma}$ such that $|E(\Gamma)|,|V(\Gamma)|>C$ can be negligible.
Since the number of diagrams with $|E|,|V|\leq C$ is finite, it suffices to analyze the diagram-wise limit. By Theorem 3.2, for all $n_{j}\gg t_{N}$, we have
@@
which completes the proof of part (ii). ∎
@@
for any fixed constant $\tau>0$, and $\|A_{N}\|_{\mathrm{op}}\leq 1+O(N^{-\frac{1}{3}}),$
then
@@
We state two standard properties of Chebyshev polynomials; their proofs can be found in the literature.
Lemma 4.6 ([EK11b, Lemma 11.2]).
@@
Under the assumption of Theorem 1.3, if $x>0$ and $x^{2}t_{N}N^{-\frac{1}{3}}<c_{0}$ for some small absolute constant, then there exist constant $c,C>0$, such that
$\mathbb{P}(\|\frac{1}{2}X\|_{\mathrm{op}}>1+xN^{-\frac{2}{3}})\leq C(1+e^{C\tau x})e^{-cx^{\frac{3}{2}}}$ (4.42)
@@
Thus use (4.41), we have for some constants $C_{1},C_{2}$,
@@
Thus, take $t_{N}\ll n=tN^{\frac{1}{3}}$, denote $\epsilon=t_{N}N^{-1/3}$ and we have
$\mathbb{E}[e^{C_{1}tN^{\frac{1}{3}}\sqrt{\|\frac{1}{2}X\|_{\mathrm{op}}-1}}]\leq C_{2}(1+e^{t\tau})(e^{C_{2}t^{\frac{3}{2}}}+e^{C_{3}\epsilon t^{2}}+(C_{4}t^{2}\epsilon)^{\frac{t}{\epsilon}}e^{C_{5}t^{3}}).$ (4.45)
By Markov inequality,
$\mathbb{P}(\|\frac{X}{2}\|_{\mathrm{op}}>1+xN^{-\frac{2}{3}})$ $\leq\mathbb{E}[e^{C_{1}tN^{\frac{1}{3}}\sqrt{(\|\frac{X}{2}\|_{\mathrm{op}}-1)_{+}}}]e^{-C_{1}t\sqrt{x}}$ (4.46)
$\leq C_{2}(1+e^{t\tau})(e^{C_{2}t^{\frac{3}{2}}}+e^{C_{3}\epsilon t^{2}}+(C_{4}t^{2}\epsilon)^{\frac{t}{\epsilon}}e^{C_{5}t^{3}})e^{-C_{1}t\sqrt{x}}.$ (by Theorem 4.4(i))
@@
$\mathbb{P}(\|\frac{X}{2}\|_{\mathrm{op}}>1+xN^{-\frac{2}{3}})\leq C_{5}(1+e^{C_{6}\tau x})(e^{-C_{7}x^{\frac{3}{2}}}+e^{C_{8}\epsilon x^{2}-C_{9}x^{\frac{3}{2}}}+(C_{10}x^{2}\epsilon)^{C_{11}x\epsilon^{-1}}e^{C_{12}x^{3}}).$ (4.47)
If $\epsilon\leq\frac{C_{9}}{2C_{8}\sqrt{x}}$ and $C_{10}x^{2}\epsilon\leq e^{-1}$ and $2C_{12}x^{3}\leq C_{11}x\epsilon^{-1}$, we have
@@
Lastly, for the third sum $\Sigma_{3}$, we need to introduce a change of variables $\lambda_{i}=2+N^{-2/3}y_{i}$ and apply Lemma 4.8. Combine these with the matching moment estimates in (4.54) and (4.55), we conclude from the continuity theorem [LZ24b, Theorem B.10] that in the supercritical regime, all $k$-point correlation functions of $X$ converge weakly to those of $\widetilde{X}$.
This indeed completes the proof by following the almost same strategy pioneered by Soshnikov [Sos99] and Sodin [Sod10]. ∎
An interesting corollary is the sinc transform of Airy process.
@@
Given a typical diagram $\Gamma$, for $t_{1},\dots,t_{s}>0$ a diagram function associated with a family of linear constraint inequalities for edge–weight variables $\{\alpha_{e}\geq 0\}$
@@
where $c_{i}(e)\in\{0,1,2\}$ counts the times of $e$ appears in $\partial D_{i}$, is defined as
@@
where $C_{\Gamma}>0$ is a combinatorial constant depending only on $\Gamma$, and for each open cycle $b_{j}\subset E_{b}(\Gamma)$
$\ell_{b_{j}}:\;=\;\sum_{e\in b_{j}}\alpha_{e}.$ (4.62)
@@
$N^{1/3}(a_{j}-1)\;\longrightarrow\;\tau_{j},\quad j=1,\dots,q.$ (4.64)

### v2

@@ 2.1 Ribbon graph expansion
The correspondence between ribbon graphs on orientable surfaces and Hermitian random matrix ensembles (e.g., GUE) is classical and well-studied; see, for example, [harer1986euler, kontsevich1992intersection, okounkov2000random, lando2004graphs]. For real ensembles (e.g., GOE), similar but subtler relations exist, involving ribbon graphs on possibly non-orientable surfaces, as formalized via the notion of Möbius graphs in [mulase2003duality].
Informally, a ribbon graph is a graph embedded into a compact surface (possibly with boundary), such that the complement of the embedding is a disjoint union of open disks, called faces. It therefore defines a cell decomposition of the surface and uniquely determines its topological type (orientable or not). Recent developments include the study of ribbon graphs with boundary, which arise naturally in open intersection theory; see, e.g., [buryak2017matrix, tessler2023combinatorial].
@@
Label the vertices in cyclic order as $v_{1},\dots,v_{k}$, and define directed edges as $\vec{e}_{j}=\overrightarrow{v_{j}v_{\gamma(j)}}$.
@@
  • In the real case ($\beta=1$), glue $\vec{e}_{s}$ and $\vec{e}_{t}$ either in the same or opposite direction (possibly yielding a non-orientable surface).
After gluing, we identify $v_{s}=v_{\gamma(t)}$ and $v_{t}=v_{\gamma(s)}$. Edges in $J^{c}$ are left unglued and correspond to the entries of external deformation $A_{N}$, which are called open edges. This construction leads to an $s$-cell ribbon graph with perimeter $(k_{1},\dots,k_{s})$ on a compact surface. For an illustration, see Figures 2 and 2.
@@
Now we introduce some definitions that have appeared in the previous articles, for instance,[geng2024outliers].
@@
A diagram $\Gamma$ is a ribbon graph satisfying the following two conditions:
@@
For the diagram, we will frequently use the following notation:
  • Edge set: let $E(\Gamma)$ denote the set of all edges of $\Gamma$, with $E_{\mathrm{int}}(\Gamma)$ and $E_{\mathrm{b}}(\Gamma)$ denoting the sets of interior and open (boundary) edges, respectively;
  • Vertex set: let $V(\Gamma)$, $V_{\mathrm{int}}(\Gamma)$ and $V_{\mathrm{b}}(\Gamma)$ denote the sets of all, interior, and boundary vertices, respectively;
  • Boundary set: for each $j\in[s]$, let $\partial D_{j}\subset E(\Gamma)$ denote the set of edges forming the boundary of face $j$.
  • Diagram set: let $\mathscr{D}_{s;\beta}$ be the set of all $s$-cell diagrams (on orientable surfaces when $\beta=2$) and $\mathscr{D}_{s;\beta}^{*}\subset\mathscr{D}_{s;\beta}$ be the subset of connected diagrams.
@@
    Record a weight $w_{e}$ on each edge $e$, which is equal to one plus the number of divalent vertices removed along it.
@@
We now state the ribbon graph expansion for the deformed Gaussian IRM ensemble as established in [geng2024outliers].
Proposition 2.4 ([geng2024outliers]).
@@
with the convention that for each trivial diagram (i.e., a single isolated vertex), the contribution is $N\delta_{l_{i},0}$. Here, the outer sum is taken over all vertex labelings $\eta:V(\Gamma)\to[N]$. Then for any integers $m_{1},\dots,m_{s}\geq 0$,
@@
This follows directly from Propositions 2.5 and 2.6 of [geng2024outliers]. ∎
@@ 2.2 Graphical expansion for Chebyshev moments
The Chebyshev polynomials of first and second kind $T_{n}(x)$ and $U_{n}(x)$ with $n\geq 0$ are defined as
@@
where the prime notation $\sum^{\prime}$ is over all $j\equiv m\pmod{2}$ and the term of $j=0$ is to be halved if there is one; see [mason2002chebyshev, eq (2.14)]. In the reverse direction the exact formula for Chebyshev polynomials in terms of powers of $x$ is
@@
Let $X$ be the deformed Gaussian matrix in Definition 1.1, then for any non-negative integers $n_{j}\geq 0$, we have a diagrammatic expansion
@@
$=\sum_{\begin{subarray}{c}1\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}\pmod{2}\end{subarray}}\sum_{\Gamma\in\mathscr{D}_{s;\beta}}\mathfrak{F}_{\Gamma}\bigl(\{n_{j}\}_{j=1}^{s}\bigr)$
@@ 3 Analysis of diagram functions
In this section, we focus on analyzing key analytic properties of diagram functions $F_{\Gamma}(\{n_{j}\})$ in (2.16) for any connected diagram $\Gamma\in\mathscr{D}_{s;\beta}^{*}$, including upper-bound estimates and precise asymptotic behavior.
@@ 3.1 Upper bound estimates
$\sum_{w_{e}\leq t_{N},e\in S\cap\mathcal{F}^{c}}\prod_{e\in S\cap\mathcal{F}^{c}}p_{w_{e}}(\eta(x),\eta(y))\leq\big(\frac{\gamma t_{N}}{N}\big)^{|S\cap\mathcal{F}^{c}|},$ (3.7)
where $|S\cap\mathcal{F}|$ denotes the number of edges. While for every edge from $S^{c}\cap\mathcal{F}^{c}$ we see from the assumption (B2) that $p_{w_{e}}(\eta(x),\eta(y))\leq(1+\delta)N^{-1}$, which implies
@@
This upper bound is irrelevant to the labeling on open vertices and weight on open edges. Since the open edges are also connected end to end into the union of several circles (as the boundary of Riemann surface are union of circles) for any given circle with $k$ open vertices, by the spectral decomposition $A_{N}=\sum_{t=1}^{r}a_{t}\mathbf{q}_{t}\mathbf{q}_{t}^{*}$ with $Q=(\mathbf{q}_{1},\ldots,\mathbf{q}_{N})$ defined in (1.5), we have
@@
If $|E|\geq{n}/{t_{N}}$, we also fix the labeling of one vertex and designate the remaining vertices as $V_{\mathrm{int}}$ with corresponding edges $E_{\mathrm{int}}$. Noting that $|V_{\mathrm{b}}|=|E_{\mathrm{b}}|=0$ and setting $a=0$, we follow the procedure in (3.18) to obtain
@@ 3.2 Asymptotics for diagram functions
$\bigg|\prod_{(x,y)\in E_{\mathrm{int}}}p_{w_{e}}(\eta(x),\eta(y))-\prod_{(x,y)\in E_{\mathrm{int}}}\frac{1}{N}\bigg|\mathbf{1}\left(\{w_{e}\leq t_{N}\}_{e\in S}\right)\cdot\mathbf{1}\left(\{w_{e}\geq t_{N}\}_{e\in S^{c}}\right).$ (3.24)
@@
Since $r$ and $|E_{\mathrm{int}}|$ are fixed, the desired result
@@ 4.1 Mixed moments in Gaussian case
The construction of $\phi$ follows the vertex-splitting procedure introduced in [feldheim2010universality] (see e.g. Step (iii) in Definition II.1.3, Figure 5 in [geng2024outliers] or Lemma 2.22 in [liu2023edge]). The key idea is to iteratively replace each unmarked vertex of degree $d>3$ with a tree of trivalent vertices connected via auxiliary edges, such that the total number of edges increases by d3d-3 and the total number of vertices increases by d3d-3. Crucially, such a replacement preserves the difference |E||V||E|-|V|.
@@
Lemma 4.3 ([feldheim2010universality, Proposition II.2.3]; [geng2024outliers, Proposition 3.7]).
@@
For the deformed Gaussian matrix $X$ in Definition 1.1 and under the assumptions (B1)-(B2) of Definition 1.2, with $n=\sum_{j=1}^{s}n_{j}$, $a=\|A_{N}\|_{\mathrm{op}}$, the following hold.
@@
    $\mathbb{E}\left[\prod_{j=1}^{s}\operatorname{Tr}\left(U_{n_{j}}\left(\frac{X}{2}\right)\right)\right]\leq(1+a^{n})(Cn)^{s}\left\{e^{C\frac{n^{\frac{3}{2}}}{\sqrt{N}}}+(1\lor\frac{t_{N}}{n})^{s}e^{C\frac{n^{2}t_{N}}{N}}+\Big((\frac{Cn^{2}t_{N}}{N})^{[\frac{n}{t_{N}}]}\land 1\Big)e^{C\frac{n^{3}}{N}}\right\}.$ (4.9)
@@
$|E|=3\ell+s,\quad|V|=2\ell+s,\quad|E_{\mathrm{int}}|-|V_{\mathrm{int}}|=|E|-|V|=\ell,$ (4.15)
where we use the fact $|E_{\mathrm{b}}|=|V_{\mathrm{b}}|$ (see [geng2024outliers, Lemma 2.9]). Rewriting $G_{\Gamma}(n):=G(n,\ell)$ that depends only on $\ell$ and $s$, we use Lemma 4.3 to further arrive at
@@
    It is also crucial to note that
@@
Since $s$ is a fixed positive integer, put (4.18)-(4.22) together and we obtain
@@
is absolutely summable. Therefore, for any sufficiently large constant $K>0$, the contribution from those diagram functions $F_{\Gamma}$ such that $|E(\Gamma)|,|V(\Gamma)|>K$ can be negligible.
Since the number of diagrams with $|E|,|V|\leq K$ is finite, it suffices to analyze the diagram-wise limit. By Theorem 3.2, for all $n_{j}\gg t_{N}$, we have
@@
which completes the proof of part (ii).
@@
for any fixed constant $\tau>0$, and $\|A_{N}\|_{\mathrm{op}}\leq 1+O(N^{-\frac{1}{3}}),$ then
@@
We state two standard properties of Chebyshev polynomials, while their proofs can be found in the literature.
Lemma 4.6 ([erdHos2011quantum, Lemma 11.2]).
@@
Under the assumption of Theorem 1.3, if $x>0$ and $x^{2}t_{N}N^{-\frac{1}{3}}<c_{0}$ for some small absolute constant, then there exist constants $c,C>0$, such that
$\mathbb{P}\Big(\|\frac{1}{2}X\|_{\mathrm{op}}>1+xN^{-\frac{2}{3}}\Big)\leq C(1+e^{C\tau x})e^{-cx^{\frac{3}{2}}}.$ (4.42)
@@
Thus using (4.41) we have for some constants $C_{1},C_{2}$,
@@
So take $t_{N}\ll n=tN^{\frac{1}{3}}$, let $\epsilon=t_{N}N^{-1/3}$ and we have
$\mathbb{E}\Big[e^{C_{1}tN^{\frac{1}{3}}\sqrt{\|\frac{1}{2}X\|_{\mathrm{op}}-1}}\Big]\leq C_{2}(1+e^{t\tau})(e^{C_{2}t^{\frac{3}{2}}}+e^{C_{3}\epsilon t^{2}}+(C_{4}t^{2}\epsilon)^{\frac{t}{\epsilon}}e^{C_{5}t^{3}}).$ (4.45)
By Markov inequality and then by Theorem 4.4(i), we have
$\mathbb{P}\Big(\|\frac{X}{2}\|_{\mathrm{op}}>1+xN^{-\frac{2}{3}}\Big)$ $\leq\mathbb{E}\Big[e^{C_{1}tN^{\frac{1}{3}}\sqrt{(\|\frac{X}{2}\|_{\mathrm{op}}-1)_{+}}}\Big]e^{-C_{1}t\sqrt{x}}$
$\leq C_{2}(1+e^{t\tau})\big(e^{C_{2}t^{\frac{3}{2}}}+e^{C_{3}\epsilon t^{2}}+(C_{4}t^{2}\epsilon)^{\frac{t}{\epsilon}}e^{C_{5}t^{3}}\big)e^{-C_{1}t\sqrt{x}}.$ (4.46)
@@
$\mathbb{P}\Big(\|\frac{X}{2}\|_{\mathrm{op}}>1+xN^{-\frac{2}{3}}\Big)\leq C_{5}(1+e^{C_{6}\tau x})(e^{-C_{7}x^{\frac{3}{2}}}+e^{C_{8}\epsilon x^{2}-C_{9}x^{\frac{3}{2}}}+(C_{10}x^{2}\epsilon)^{C_{11}x\epsilon^{-1}}e^{C_{12}x^{3}}).$ (4.47)
When $\epsilon\leq\frac{C_{9}}{2C_{8}\sqrt{x}}$ and $C_{10}x^{2}\epsilon\leq e^{-1}$ and $2C_{12}x^{3}\leq C_{11}x\epsilon^{-1}$, we have
@@
Lastly, for the third sum $\Sigma_{3}$, we need to introduce a change of variables $\lambda_{i}=2+N^{-2/3}y_{i}$ and apply Lemma 4.8. Combine these with the matching moment estimates in (4.54) and (4.55), we conclude from the continuity theorem [liu2023edge, Theorem B.10] that in the supercritical regime, all $k$-point correlation functions of $X$ converge weakly to those of $\widetilde{X}$.
This indeed completes the proof by following the almost same strategy pioneered by Soshnikov [soshnikov1999universality] and Sodin [sodin2010spectral]. ∎
A related and interesting corollary is the evaluation of the sinc transform for correlation functions of (deformed) Airy point processes, including both real and complex cases. In the non-deformed setting, this result was first obtained by Sodin [sodin2010spectral].
@@
Given a typical diagram $\Gamma$, for $t_{1},\dots,t_{s}>0$ introduce a family of linear constraint inequalities for edge–weighted variables $\{\alpha_{e}\geq 0\}$
@@
with $c_{i}(e)\in\{0,1,2\}$ counting the times of $e$ appearing in $\partial D_{i}$, a diagram function is defined as
@@
where for each open cycle $b_{j}\subset E_{b}(\Gamma)$
$\ell_{b_{j}}:\;=\;\sum_{e\in b_{j}}\alpha_{e},$ (4.62)
and $C_{\Gamma}>0$ is a combinatorial constant depending only on $\Gamma$ defined by
$C_{\Gamma}=\lim_{n_{j}\rightarrow\infty}\frac{\#\{(w_{e})_{e\in E(\Gamma)}|t_{j},w_{e}\in\mathbb{N},~t_{j}\geq 0,\;w_{e}\geq 1,~2t_{j}+\sum\limits_{e\in\partial D_{j}}w_{e}=n_{j},~j=1,\dots,s\}}{\mathrm{Vol}\{(\alpha_{e})_{e\in E(\Gamma)}|\alpha_{e}\in\mathbb{R}^{+},\sum_{e\in E(\Gamma)}c_{i}(e)\,\alpha_{e}\;\leq\;n_{j},\quad j=1,\dots,s\}}.$ (4.63)
@@
$N^{1/3}(a_{j}-1)\;\longrightarrow\;\tau_{j},\quad j=1,\dots,q,$ (4.65)